OpenStudy (anonymous):

a. write the expression in radical notation will show it on the equation bar b. evaluate the radical expression

OpenStudy (anonymous):

\[8^{-2/3}\]

OpenStudy (anonymous):

when written it would it look like this\[\sqrt[3]{8^{-2}}\]

OpenStudy (mathstudent55):

correct so far

OpenStudy (mathstudent55):

Be careful, the exponent of 8 is -2 not 2

OpenStudy (mathstudent55):

8^2 = 64 8^-2 = 1/8^2 = 1/64

OpenStudy (anonymous):

so the answer would be 1/64 for the evaluate rational expression

OpenStudy (mathstudent55):

it's cubic root of 1/64, which can be simplified

OpenStudy (mathstudent55):